OpenStudy (pokemon23):

Solve each equation for x: 4x+ 9-2x=31 show all work Saifoo once again does the job right :) Solve each equation for x: 4x+ 9-2x=31 show all work Saifoo once again does the job right :) @Mathematics

hero (hero):

2x + 9 = 31 2x = 22 x = 11

OpenStudy (pokemon23):

ty hero :0

OpenStudy (lilg132):

4x -2x = 31 -9 2x = 22 x = 11

OpenStudy (pokemon23):

and lilg

OpenStudy (saifoo.khan):

lol, im late this time.